OpenStudy (anonymous):

At a market, a customer who purchases 3lb of potato salad and 4lb of coleslaw pays a total of $10.75. Another customer pays a total of $4.75 to buy 1 lb of potato salad and 2 lb of coleslaw. What is the cost to purchase 2 lb of potato salad and 3lb of coleslaw?

OpenStudy (anonymous):

solve \[3x+4y=10.75\] \[x+2y=4.75\] get \[x=1.25,y=1.75\]

OpenStudy (anonymous):

then compute \[2\times 1.25+3\times 1.75\] for your answer

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

yw